OVERTONE AND COMBINATION BAND SPECTROSCOPY OF H_{3}^{+}

\underline{BENJAMIN J. McCALL^{a}}, Department of Chemistry and Department of Astronomy and Astrophysics, The University of Chicago, Chicago, IL 60637; THERESE R. HUET, Laboratoire de Spectroscopie Hertzienne, Centre d’Etudes et de Recherches Lasers et Applications, Universite de Lille, 59 655 Villenueve d’Ascq cedex, France; JAMES K. G. WATSON, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, Ontario, Canada K1A 0R6; TAKESHI OKA, Department of Chemistry and Department of Astronomy and Astrophysics, The University of Chicago, Chicago, IL 60637.

We report our detection of H_{3}^{+} rovibrational transitions between 7700 and 8200 cm^{−1}. This work employed a New Focus external cavity tunable diode laser, a positive column discharge, and the velocity modulation technique for enhanced sensitivity to molecular ions. This frequency range encompasses transitions of the 3ν_{2}^{1} ← 0 and ν_{1} + 2ν_{2} ← 0 bands — the highest energy transitions studied to date.

As the simplest three-atom molecule, H_{3}^{+} has long served as a benchmark for ab initio calculations. We have compared our observed frequencies with our predictions, and expect that this work (along with detection of even higher overtones) will lead to better ab initio treatments of this important molecule.

aSupported by the Fannie and John Hertz Foundation